OpenStudy (anonymous):

Solve: x over 21 is less than or equal to negative 6 A. x is less than or equal to negative 126 B. x is greater than or equal to negative 3.5 C. x is less than or equal to negative 3.5

terenzreignz (terenzreignz):

Multiplication property works in a very similar fashion for inequalities as it does for equality, as long as what you multiply to both sides is positive... \[\huge \dfrac{x}{21}\le-6\]

terenzreignz (terenzreignz):

So you can multiply both sides of the inequality by 21... \[\huge 21\cdot\frac{x}{21}\le -6\cdot 21\]\[\huge \cancel{21}\cdot\frac{x}{\cancel{21}}\le -6\cdot 21\]

OpenStudy (anonymous):

-6.5 isn't a answer though.

terenzreignz (terenzreignz):

How would you get -6.5?

OpenStudy (anonymous):

Meant -6.21 Sorry

terenzreignz (terenzreignz):

Oh, sorry, that's not a decimal point... \[\huge \cancel{21}\color{red}{\times}\frac{x}{\cancel{21}}\le -6\color{red}{\times} 21\]

terenzreignz (terenzreignz):

The raised dot is a symbol for multiplication... you'd do well to remember that :)

OpenStudy (anonymous):

Ok Thank you I get it now!

terenzreignz (terenzreignz):

No problem :)